OpenStudy (anonymous):

Evaluate: 3^2 + (8 − 2) • 4 − 6/3. 30 58 28 31

OpenStudy (anonymous):

I got 9..? But thats not an option?

OpenStudy (anonymous):

okai so lets see, 9 +6 x 4 - 2 15 x 2 30

OpenStudy (anonymous):

See what I did?

OpenStudy (anonymous):

remember you should factor down the fraction as much as possible so 6/3 = 2/1 which is 2

OpenStudy (anonymous):

no im lost. What i did was 9+24-2 and it cam out as 34-2= 32

OpenStudy (anonymous):

Ok haha. i meant *31

OpenStudy (anonymous):

okai let me explain then. Take it one step at a time first start with the first step which is 3^2 this equals 9. So write down 9. Next you just write the symbol inbetween them them which is a +. Okai now you do what is in the parenthese which is 8-2 easy enough that is 6 so write 6. then have times 4, and you're going to leave that alone for right now. so just go ahead and write that in the new equation. Then we do the fraction, 6/3 = 2/1 = 2 so write the symbol - and write 2 so now the new equation should be 9+6 x 4-2

OpenStudy (anonymous):

Now you just do 9+6 which is 15 and do 4-2 which is 2 and leave the x inbetween. so now we have 15 x 2 and the final answer is 30.

OpenStudy (anonymous):

Understand now?

OpenStudy (anonymous):

The order of operations is 1. Parentheses 2. Exponents 3. Multiplication or Division, left to right 4. Addition or Subtraction, left to right Multiplication and division are equal in order and should be performed from left to right. Addition and subtraction are equal in order and should be performed from from left to right.

OpenStudy (anonymous):

in the other one you said over 3 lol in this case it is 31

OpenStudy (anonymous):

The entire equation is not over 3 so no when you factor down 6/3 it comes out to be 2/1 which is 2. Therefore 30 is the correct answer.

OpenStudy (anonymous):

oh okay i tthinkkk i understand now.

OpenStudy (anonymous):

I hope c: Like what that other guy said use the order of operations which is what I did. I just tried to break it down for you as much as I could.